Planning week 17 - 24 December 2016

Everything in this plan was lost except studies on 17 December up to 18:01. The problem is believed to be due to a single event upset (SEU) which caused an error in the MDP. I've left the notes I originally wrote for reference.

The main thing going on this week was a set of coordinated campaigns with ALMA, taking place between 13 and 20 UT on Sat-Sun and Tue-Thu. HOP 79 (irradiance scan) was supported on 20-21 Dec.

The week began with very low GOES activity, and a large north coronal hole rotating around the east limb. An unnumbered AR was sitting in a large island within this coronal hole, and this was the main target for OP1. A bright point rapidly expanded to a small AR (12619) on 18 December and became the default target for OP2. Another bright point emerged on 19 December and was unusual in having an "anti-Hale" magnetic configuration. This was named AR 12620 and became the default target for OP3.

I wanted to take advantage of the large coronal hole for jet or plume studies, but I saw very little jet activity and very few plumes, and the only study I ran was on Mon-Tue.

Saturday 17 Dec (OP 1)

AIA 193 image

As XRT was mostly in bake-out for this OP I got 34% of the DR allocation (4186 Mbits).

I ran Atlas_60 on QS near disk center at 11:10, then HOP 328 (with ALMA) was run between 12:30 and 15:30. The target was spicules at the east limb, and I just ran Cool_loop_response continuously.

The unnumbered AR had a very bright fan loop that was next to a lane of coronal hole, so I decided to run my old study PRY_footpoints_HI2 six times, spread over a 48-hour period. The times are:

17/16:16 -> with IRIS (UPDATE: this data-set was successful, all later ones were lost)

The raster should include the fan loop and the coronal hole. **UPDATE: the fan loop was significantly fainter by the time the observation started, and continued to fade. The AR was essentially dead by the end of the OP.**

Sunday 18 Dec (OP1)

AIA 193 image

The PRY_footpoint_HI2 raster was run three times (see above).

The ALMA campaign was supported again between 18:30 and 20:00 with the target a prominence at the west limb. I thought the best study was simply an Atlas_60, which I scheduled at 18:32.

Monday 19 Dec (OP1)

AIA 193 image

No ALMA campaign today. The PRY_footpoints_HI2 was run twice (see above).

From 10:21 to 15:28 I scheduled Cool_loop_stare on the bright fan loop of the AR. I also followed it with a single run of Cool_loop_resp_v2 for some context. **This was coordinated with IRIS.** **UPDATE: the fan loop was extremely weak by this point.**

For the period 18:15 to 09:07 (20th) I scheduled Cool_loop_narrow for monitoring small-scale heating and dynamics in a patch of coronal hole near disk center.**This was coordinated with IRIS.** **This raster ended up being almost directly south of AR 12619.**

Tuesday 20 Dec (OP1)

The coronal hole observation continued into the 20th.

Tuesday 20 Dec (OP2)

AIA 193 image

The small, non-flaring AR 12619 was the default target for this OP. I began with Atlas_60 at 10:53 UT, and then ran ar_evolution from 12:00 UT for 5.5 hours. This included the period for HOP 326, an AR plage study coordinated with IRIS and ALMA.

After some engineering studies, I resumed with ar_evolution from 18:20 UT for 3 hours, in support of HOP 328 (generic ALMA AR observation, with IRIS).

The 20 pointings for HOP 79 (north-south scan) began at 21:19 UT and continued to 21-Dec.

Wednesday 21 Dec (OP2)

AIA 193 image

HOP 79 continued until 11:42 UT. At 12:42 I resumed AR tracking with ar_evolution for 4 hours, as part of HOP 328 again. An AR spectral atlas was performed at 16:46 UT, and then I resumed AR tracking for the rest with ar_evolution at 18:02 (12 hours; 46 rasters).

Thursday 22 Dec (OP2)

AIA 193 image

AR 12619 tracking continued, with an additional sequence at 09:20 UT (3 hours; 12 rasters).

Thursday 22 Dec (OP 3)

AIA 193 image

Since AR 12620 was near the limb, SOT gave up a lot of the telemetry, and I got 39.%.

For this OP I tried to run the study ar_evolution on AR 12620 most of the time.

Prior to HOP 329 (ALMA) I ran ar_evolution at 11:25 (3 repeats). HOP 329 was supported on AR 12620 with the sit-and-stare study SI_Mercury_slit for four hours from 12:45. Note this study has a very high data rate.

HOP 328 (generic ALMA campaign was supported on AR 12620 with 9 repeats of ar_evolution from 17:52. AR tracking of AR 12620 then began with 34 repeats of ar_evolution from 20:30.

Friday 23 Dec (OP 3)

AIA 193 image

The ar_evolution sequence on AR 12620 continued until 05:23 and then I followed with Atlas_60.

The next ar_evolution sequence (37 rasters) began at 06:37, and then another one of 37 rasters began at 18:22.

Saturday 24 Dec (OP 3)

AIA 193 image

The ar_evolution sequence on AR 12620 continued until 04:03 and then I followed with Atlas_60.

The final ar_evolution sequence of 10 rasters began at 06:09.

Page maintained by Dr Peter R Young.